OpenStudy (anonymous):

Order the following from least to greatest; 6.5, 6/5, 600%, 1/6, -60, 1/5 Please show work I dont know how to turn fractions and percents into decimals

OpenStudy (anonymous):

To turn fractions into decimals, you'll want to divide the numerator by the denominator, so 6/5= 1.2, and 1/6=.167, and 1/5=.2. To make a percent a decimal, just multiply it by .01 or 1/100. So 600%=6. 1 Therefore the order will go in -60, 1/6, 1/5, 6/5, and 600%